# Comparing Hubbard parameters from linear-response theory and Hartree-Fock-based approach

### v1

@@ INTRODUCTION
Density-functional theory (DFT) Hohenberg:1964, Kohn:1965 underpins modern first-principles simulations in physics, chemistry, and materials science. Despite its remarkable success, progress hinges on the development of increasingly accurate exchange-correlation (xc) functionals. Standard approximations, such as the local spin-density approximation (LSDA) and spin-polarized generalized-gradient approximation ($\sigma$-GGA), are invaluable for many systems, yet they perform poorly for transition-metal (TM) and rare-earth (RE) compounds, where localized $d$ and $f$ states suffer from large self-interaction errors (SIEs) Perdew:1981, MoriSanchez:2006. To address these limitations, several beyond-standard approaches have been introduced. Among the most prominent are Hubbard-corrected DFT (DFT+$U$ anisimov:1991, Liechtenstein:1995, Dudarev:1998 and its extension DFT+UU+$V$ Campo:2010), meta-GGA functionals (e.g., SCAN Sun:2015 and its variants Bartok:2019, Furness:2020), and hybrid functionals (e.g., PBE0 Adamo:1999 and HSE06 Heyd:2003, Heyd:2006). Combinations such as SCAN+$U$ have also been explored Gautam:2018, Long:2020, Kaczkowski:2021, Artrith:2022. Each strategy embodies a distinct philosophy: DFT+$U$ selectively corrects partially filled orbitals (typically $d$ and $f$) Kulik:2006, Kulik:2008; meta-GGA incorporates the kinetic-energy density and satisfies a broad set of exact constraints (17 in the case of SCAN); hybrid functionals mix a fraction of Fock exchange with (semi-)local functionals. While these methods improve predictive power, they each come with intrinsic trade-offs and limitations Timrov:2022b. Among them, Hubbard-corrected DFT has become especially popular for its ability to enhance accuracy at modest computational cost Himmetoglu:2014, provided reliable Hubbard parameters are available. These can be obtained either semi-empirically (by fitting to experimental data or higher-level calculations such as hybrids or $GW$ Hedin:1965) or through systematic calibration. Examples include tuning $U$ to reproduce binary metal formation energies Wang:2006, Kubaschewski:1993 or to match properties such as band gaps, magnetic moments, lattice parameters, and enthalpies LeBacq:2004, Aykol:2014, Isaacs:2017.
@@
While semi-empirical fitting of Hubbard parameters can be effective, it becomes impractical in materials discovery where experimental data are unavailable or when multiple target properties must be simultaneously reproduced, sometimes requiring advanced optimization schemes such as Bayesian approaches Yu:2020, Tavadze:2021. A natural alternative is to compute Hubbard parameters from first principles using methods such as constrained DFT (cDFT) Dederichs:1984, Mcmahan:1988, Gunnarsson:1989, Hybertsen:1989, Gunnarsson:1990, Pickett:1998, Solovyev:2005, Nakamura:2006, Shishkin:2016, Hartree-Fock-based schemes Mosey:2007, Mosey:2008, Andriotis:2010, Agapito:2015, Lee:2020, TancogneDejean:2020, or the constrained random-phase approximation (cRPA) Springer:1998, Kotani:2000, Aryasetiawan:2004, Aryasetiawan:2006. However, predicted values can differ substantially among these approaches Himmetoglu:2014, Aryasetiawan:2006, Liu:2023, and systematic comparisons of their theoretical foundations remain scarce, although notable progress is beginning to emerge Carta:2025. Moreover, the dependence of $U$ on factors such as Hubbard projector choice, pseudopotentials, oxidation state, and base xc functional further complicates reproducibility Kulik:2008, Shishkin:2016, Carta:2025, Mahajan:2021, Hubparam:2023. Therefore, further work similar to Ref. Carta:2025 is needed to systematically compare and benchmark different first-principles approaches for computing Hubbard parameters.
@@
Here, we present a comparative study of two such first-principles approaches. The first is based on cDFT formulated within linear-response theory (LRT) Cococcioni:2005, which determines on-site $U$ and inter-site $V$ by enforcing piecewise linearity of the total energy with respect to the occupation of localized orbitals (Hubbard manifold), thereby reducing SIEs Himmetoglu:2011 and improving predictive accuracy. Standard LRT requires supercell calculations, but its reformulation within density-functional perturbation theory (DFPT) Timrov:2018, Timrov:2021 replaces these with monochromatic reciprocal-space perturbations in a unit cell. This substantially reduces computational overhead and has enabled applications to increasingly complex systems Mahajan:2022, Timrov:2023, Binci:2023, Gebreyesus:2023, Haddadi:2024, Bonfa:2024, Macke:2024, Warda:2025. The second approach for computing $U$ is the Hartree-Fock-based Agapito, Curtarolo, and Buongiorno Nardelli (ACBN0) functional Agapito:2015 and its extension to inter-site $V$ dubbed as eACBN0 Lee:2020, TancogneDejean:2020, which derives Hubbard parameters directly within self-consistent-field (SCF) cycles from an approximate form of the screened Hartree-Fock interaction energy. Using projections of Kohn-Sham (KS) states onto localized orbitals, $U$ and $V$ are updated at each iteration of the SCF cycle with negligible additional cost relative to standard DFT. This method has been shown to capture electronic structures and dynamical properties with good accuracy Yang:2021, Yang:2022, and it holds promise for accurate evaluations of electron-phonon interactions Jang:2023, Yang:2025.
@@
The DFT+UU+$V$ approach. In DFT+UU+$V$ Campo:2010, Himmetoglu:2014, a corrective Hubbard energy $E_{U+V}$ is added to the standard DFT energy $E_{\mathrm{DFT}}$:
@@
Unlike DFT+$U$, which only accounts for on-site interactions scaled by $U$, DFT+UU+$V$ also includes inter-site interactions between an atom and its neighbors, scaled by $V$. In the simplified rotationally-invariant formulation Dudarev:1998, the extended Hubbard correction reads Timrov:2021:
@@
$n^{IJ\sigma}_{mm^{\prime}}=\sum_{v,\mathbf{k}}f^{\sigma}_{v,\mathbf{k}}\langle\psi^{\sigma}_{v,\mathbf{k}}|\phi^{J}_{m^{\prime}}\rangle\langle\phi^{I}_{m}|\psi^{\sigma}_{v,\mathbf{k}}\rangle\,$, (3)
where $v$ and $\sigma$ label bands and spin, $\mathbf{k}$ runs over the first Brillouin zone (BZ), and $f^{\sigma}_{v,\mathbf{k}}$ are the KS occupations. In Eq. (3), $m\in I$ and $m^{\prime}\in J$, meaning that $m$ runs over the magnetic quantum numbers associated with a fixed principal quantum number $n$ and orbital angular momentum $l$ of atom $I$, and analogously $m^{\prime}$ for atom $J$. The corresponding generalized KS equation can be written as Campo:2010,
@@
where $\hat{\mathcal{H}}_{\mathrm{KS}}^{\sigma}$ is the KS Hamiltonian of the standard DFT, $\varepsilon_{v,{\bf k}}^{\sigma}$ is the $v$-th band energy at momentum $\bf k$} and spin $\sigma$, and $\hat{\mathcal{V}}_{U+V}^{\sigma}$ is the Hubbard potential that is derived from the Hubbard energy (2) and which reads Timrov:2021:
@@
It is important to note that the computed values of $U^{I}$ and $V^{IJ}$ depend on the choice of Hubbard projector functions $\phi^{I}_{m}(\mathbf{r})$ used to define the occupation matrix in Eq. (3) and the Hubbard potential in Eq. (5) Lee:2020, KirchnerHall:2021, Mahajan:2021. In the following, we briefly describe how Hubbard parameters are defined and computed within LRT and eACBN0. For both methods, we employ the same projector functions, namely atomic orbitals orthogonalized using the Löwdin method Lowdin:1950, Mayer:2002, to ensure a fair and consistent comparison of the computed $U^{I}$ and $V^{IJ}$ values.
Hubbard parameters from LRT. Within the LRT framework, Hubbard parameters are defined from the piecewise linearity condition of the total energy [Eq. (1)] with respect to the occupation of the Hubbard manifold and computed as Cococcioni:2005, Campo:2010:
@@
DFPT allows this calculation to be performed without supercells, by reformulating LRT in terms of monochromatic reciprocal-space perturbations in the unit cell. This significantly reduces computational cost Timrov:2018, Timrov:2021. The response of the occupation matrix is then computed as:
@@
where $\mathbf{q}$ is the wavevector of the monochromatic perturbation, $N_{\mathbf{q}}$ is the total number of perturbations, and $\Delta_{\mathbf{q}}^{s^{\prime}}n^{s\sigma}_{mm^{\prime}}$ is the lattice-periodic response of the occupation matrices to a $\mathbf{q}$-specific perturbation. Here, $I\equiv(l,s)$ and $J\equiv(l^{\prime},s^{\prime})$, with $s$ and $s^{\prime}$ indexing atoms within the unit cell and $l$ and $l^{\prime}$ indexing unit cells; $\mathbf{R}_{l}$ and $\mathbf{R}_{l^{\prime}}$ are the corresponding Bravais lattice vectors. The quantities $\Delta{\mathbf{q}}^{s^{\prime}}n^{s\sigma}_{mm^{\prime}}$ are obtained from the response KS wavefunctions, which are computed by solving $\mathbf{q}$-specific Sternheimer equations. Further details on the DFPT implementation can be found in Refs. Timrov:2018, Timrov:2021. The efficiency of DFPT for calculating $U^{I}$ and $V^{IJ}$ has been demonstrated in numerous studies, with several recent highlights Gelin:2024, Uhrin:2025, Bastonero:2025, Chang:2025, Binci:2025, dosSantos:2025.
Hubbard parameters from eACBN0. In the eACBN0 method, the Hubbard parameters are defined by equating the corrective Hubbard energy [Eq. (2)] with the Hartree-Fock interaction energy Agapito:2015, Lee:2020. This leads to the following expressions:
@@
where $m\in I$ and $m^{\prime}\in J$, and the summations $\sum_{\{I\}}$ and $\sum_{\{J\}}$ mean that we sum up contributions coming from all equivalent atoms of the same type. For the case of $U^{I}$-only calculations ($I=J$), the second term in Eq. (13) is dropped Lee:2020, so $\bar{N}^{II,\sigma}_{v,\mathbf{k}}$ reduces to the Mulliken charge of atom $I$ in state $\psi^{\sigma}_{v,\mathbf{k}}(\mathbf{r})$. Moreover, $\bar{N}^{IJ,\sigma}_{v,\mathbf{k}}$ can be viewed as a position-dependent mixing parameter within the Hartree-Fock approximation, encoding on-site and inter-site screening effects in the localized subspace. It is important to note that for a fixed ground-state density, the $U^{I}$ values computed in ACBN0 remain unchanged upon inclusion of inter-site $V^{IJ}$ terms in eACBN0. However, in self-consistent implementations of eACBN0, a coupling between $U^{I}$ and $V^{IJ}$ emerges because the ground state is obtained under the influence of both on-site and inter-site terms in the Hubbard potential (5), with self-consistency enforced for both the electron density and the Hubbard parameters. Further details on the eACBN0 formalism can be found in Refs. Agapito:2015, Lee:2020, TancogneDejean:2020.
@@ RESULTS
Figure 1: Protocol for computing Hubbard parameters within the “self-consistent” scheme using (a) LRT (DFPT) Timrov:2021 and (b) eACBN0 Lee:2020. In the LRT workflow, Hubbard parameters are evaluated self-consistently after the electronic charge density has been fully converged. In contrast, the eACBN0 scheme updates the charge density and the Hubbard parameters simultaneously until convergence is reached. In both workflows, atomic positions are fixed and no structural relaxations are performed.
Hubbard parameters can be computed at different levels of self-consistency. The simplest approach is a “one-shot” calculation, where the parameters are evaluated using the converged DFT charge density while keeping the geometry fixed. A more advanced “self-consistent” approach iteratively updates both the charge density and the corresponding $U$ and $V$ values, but still at a fixed geometry. Finally, in the “structurally self-consistent” scheme, structural relaxations (atomic positions and unit-cell parameters) are included in the self-consistent cycle Timrov:2021, Bastonero:2025. In the main text, we report results obtained with the “self-consistent” scheme, i.e. calculations in which the electronic degrees of freedom are converged together with the Hubbard parameters while the geometry remains fixed. In this approach, all $U$ and $V$ values are determined consistently with the self-consistent charge density Timrov:2021, Lee:2020, as illustrated in Fig. 1. This allows us to isolate purely electronic effects from structural ones and to assess the influence of extended Hubbard interactions in LRT and eACBN0 with high internal consistency. For completeness, in the Supplemental Information (SI) we also present results obtained using the “structurally self-consistent” scheme, in which the self-consistent cycle includes full structural optimization with updated Hubbard parameters and charge density (see Fig. S1 in the SI). In this case, the crystal structure, the Hubbard parameters, and the charge density are iterated to convergence, yielding mutually consistent electronic and crystal structures. All computational details are provided in the Methods section.
@@
Since the band gap is central to this study, it is useful to recall that DFT is not a spectral theory. Local and semilocal xc functionals (LSDA, $\sigma$-GGA) notoriously underestimate band gaps in transition-metal oxides (TMOs) due to strong SIEs, particularly on TM ions. DFT+$U$ can alleviate this issue, but only when the correction is applied to states that define the gap edges (for insulators) or to states near the Fermi level (for spurious DFT metals) KirchnerHall:2021. The inter-site $V$ correction can further improve band gaps in covalent systems, but again only when applied to orbitals directly involved in gap formation. Prior work supports these points: Ref. KirchnerHall:2021 showed that DFT+$U$ with LRT-derived $U$ improves agreement with experimental band gaps compared to (semi-)local functionals, while DFT+UU+$V$ with LRT-derived parameters yields even better accuracy in TMOs with strong covalency Mahajan:2021, Mahajan:2022, Timrov:2023, Gebreyesus:2023. For $p$-block semiconductors without TM ions, Ref. Lee:2020 demonstrated that DFT+UU+$V$ with eACBN0-derived parameters accurately reproduces experimental band gaps. These insights provide the foundation for our analysis below.
@@
Figure 2: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for TMOs without on-site corrections on O-2pp states ($U_{O}$), using the workflow described in Fig. 1. (a) On-site Hubbard $U_{TM}$ parameter for 3dd orbitals of TM ions and inter-site Hubbard $V_{TM,O}$ parameter between 3dd orbitals of TM ions and 2pp orbitals of oxygen, (b) band gap, and (c) magnetic moment. LRT and eACBN0 denote the method used to obtain the Hubbard parameters. The experimental band gaps and magnetic moments for MnO Kurmaev2008prb, Elp15301991prb, Cheetham1983prb, Fender1968JCP, FeO Bowen1975JSSC, Roth1958prb, Battle1979JPCSSP, CoO Elp60901991prb, Khan1970prb, Jauch2001prb, and NiO Kurmaev2008prb, Sawatzky1984prl, Cheetham1983prb, Fender1968JCP are indicated by the dashed blue lines in panels (b) and (c), respectively.
@@
To compare the two methods under as equivalent conditions as possible, we perform LRT and eACBN0 calculations using the “self-consistent” scheme described in Fig. 1 and report the Hubbard parameters in Figs. 2 and 3. We focus first on the $U$ values shown as bar plots (unhatched) in Fig. 2 obtained in the framework of the DFT+$U$ method. A clear trend emerges where $U_{TM}$ increases and magnetic moment decreases with the number of 3dd electrons. Furthermore, both calculation methods do not alter the $U_{TM}$ values significantly when the inter-site interactions are included (compare the hatched and not hatched patterns in Fig. 2(a)). Focusing on MnO, the dominant Mn-$4s$ character of the lowest conduction band turns out to be insensitive to the Hubbard $U$ parameter applied to Mn-$3d$ states. The primary role of $U_{TM}$ is to enhance the splitting between occupied and unoccupied Mn-3dd states. This is confirmed by the partial density of states (PDOS), which shows that with increasing $U_{TM}$, occupied $d$ states shift to lower energies, while the unoccupied $d$ states (lying above the conduction band minimum) shift to higher energies Yang:2022. This mechanism is found to be general for other late TMOs like FeO, CoO, and NiO, which also exhibit a prominent TM-4ss character at the conduction band minimum at the $\Gamma$ point. For these materials, we consistently observe that while the energy splitting between occupied and unoccupied $d$ states scales proportionally with $U_{TM}$, the variation in the fundamental band gap remains minor in comparison.
@@
When the inter-site Hubbard interactions ($V_{TM,O}$) between TM 3dd and oxygen 2pp states are included, the differences between the two methods remain almost unchanged. In Fig. 2, the results with the $V_{TM,O}$ term are illustrated with hatched patterns. LRT yields relatively small $V_{TM,O}$ values (below 1 eV), resulting in negligible differences in band gaps and magnetic moments whether $V_{TM,O}$ term is included or not in the total energy calculation. In contrast, eACBN0 consistently produces larger $V_{TM,O}$ values (close to 3 eV) across all TMOs, resulting in more noticeable changes in band gaps compared to LRT, while the magnetic moments change by a small margin. This effect arises because the $V_{TM,O}$ term facilitates electron delocalization into the overlapping regions between atomic pairs. This counteracts the strong localization favored by the $U_{TM}$ term and thus reduces the local magnetic moment, as previously shown through charge density difference analysis in our earlier studies Yang:2022.

### v2

@@ INTRODUCTION
Density-functional theory (DFT) [Hohenberg:1964, Kohn:1965] underpins modern first-principles simulations in physics, chemistry, and materials science. Despite its remarkable success, progress hinges on the development of increasingly accurate exchange-correlation (xc) functionals. Standard approximations, such as the local spin-density approximation (LSDA) and spin-polarized generalized-gradient approximation ($\sigma$-GGA), are invaluable for many systems, yet they perform poorly for transition-metal (TM) and rare-earth (RE) compounds, where localized $d$ and $f$ states suffer from large self-interaction errors (SIEs) [Perdew:1981, MoriSanchez:2006]. To address these limitations, several beyond-standard approaches have been introduced. Among the most prominent are Hubbard-corrected DFT (DFT+$U$ [anisimov:1991, Liechtenstein:1995, Dudarev:1998] and its extension DFT+UU+$V$ [Campo:2010]), meta-GGA functionals (e.g., SCAN [Sun:2015] and its variants [Bartok:2019, Furness:2020]), and hybrid functionals (e.g., PBE0 [Adamo:1999] and HSE06 [Heyd:2003, Heyd:2006]). Combinations such as SCAN+$U$ have also been explored [Gautam:2018, Long:2020, Kaczkowski:2021, Artrith:2022]. Each strategy embodies a distinct philosophy: DFT+$U$ selectively corrects partially filled orbitals (typically $d$ and ff[Kulik:2006, Kulik:2008]; meta-GGA incorporates the kinetic-energy density and satisfies a broad set of exact constraints (17 in the case of SCAN); hybrid functionals mix a fraction of Fock exchange with (semi-)local functionals. While these methods improve predictive power, they each come with intrinsic trade-offs and limitations [Timrov:2022b]. Among them, Hubbard-corrected DFT has become especially popular for its ability to enhance accuracy at modest computational cost [Himmetoglu:2014], provided reliable Hubbard parameters are available. These can be obtained either semi-empirically (by fitting to experimental data or higher-level calculations such as hybrids or $GW$ [Hedin:1965]) or through systematic calibration. Examples include tuning $U$ to reproduce binary metal formation energies [Wang:2006, Kubaschewski:1993] or to match properties such as band gaps, magnetic moments, lattice parameters, and enthalpies [LeBacq:2004, Aykol:2014, Isaacs:2017].
@@
While semi-empirical fitting of Hubbard parameters can be effective, it becomes impractical in materials discovery where experimental data are unavailable or when multiple target properties must be simultaneously reproduced, sometimes requiring advanced optimization schemes such as Bayesian approaches [Yu:2020, Tavadze:2021]. A natural alternative is to compute Hubbard parameters from first principles using methods such as constrained DFT (cDFT) [Dederichs:1984, Mcmahan:1988, Gunnarsson:1989, Hybertsen:1989, Gunnarsson:1990, Pickett:1998, Solovyev:2005, Nakamura:2006, Shishkin:2016], Hartree-Fock-based schemes [Mosey:2007, Mosey:2008, Andriotis:2010, Agapito:2015, Lee:2020, TancogneDejean:2020], or the constrained random-phase approximation (cRPA) [Springer:1998, Kotani:2000, Aryasetiawan:2004, Aryasetiawan:2006]. However, predicted values can differ substantially among these approaches [Himmetoglu:2014, Aryasetiawan:2006, Liu:2023], and systematic comparisons of their theoretical foundations remain scarce, although notable progress is beginning to emerge [Carta:2025]. Moreover, the dependence of $U$ on factors such as Hubbard projector choice, pseudopotentials, oxidation state, and base xc functional further complicates reproducibility [Kulik:2008, Shishkin:2016, Carta:2025, Mahajan:2021, Hubparam:2023]. Therefore, further work similar to Ref. Carta:2025 is needed to systematically compare and benchmark different first-principles approaches for computing Hubbard parameters.
@@
Here, we present a comparative study of two such first-principles approaches. The first is based on cDFT formulated within linear-response theory (LRT) [Cococcioni:2005], which determines on-site $U$ and inter-site $V$ by enforcing piecewise linearity of the total energy with respect to the occupation of localized orbitals (Hubbard manifold), thereby reducing SIEs [Himmetoglu:2011] and improving predictive accuracy. Standard LRT requires supercell calculations, but its reformulation within density-functional perturbation theory (DFPT) [Timrov:2018, Timrov:2021] replaces these with monochromatic reciprocal-space perturbations in a unit cell. This substantially reduces computational overhead and has enabled applications to increasingly complex systems [Mahajan:2022, Timrov:2023, Binci:2023, Gebreyesus:2023, Haddadi:2024, Bonfa:2024, Macke:2024, Warda:2025]. The second approach for computing $U$ is the Hartree-Fock-based Agapito, Curtarolo, and Buongiorno Nardelli (ACBN0) functional [Agapito:2015] and its extension to inter-site $V$ dubbed as eACBN0 [Lee:2020, TancogneDejean:2020], which derives Hubbard parameters directly within self-consistent-field (SCF) cycles from an approximate form of the screened Hartree-Fock interaction energy. Using projections of Kohn-Sham (KS) states onto localized orbitals, $U$ and $V$ are updated at each iteration of the SCF cycle with negligible additional cost relative to standard DFT. This method has been shown to capture electronic structures and dynamical properties with good accuracy [Yang:2021, Yang:2022], and it holds promise for accurate evaluations of electron-phonon interactions [Jang:2023, Yang:2025].
The DFT+U+V approach.
@@
In DFT+UU+$V$ [Campo:2010, Himmetoglu:2014], a corrective Hubbard energy $E_{U+V}$ is added to the standard DFT energy $E_{\mathrm{DFT}}$:
@@
Unlike DFT+$U$, which only accounts for on-site interactions scaled by $U$, DFT+UU+$V$ also includes inter-site interactions between an atom and its neighbors, scaled by $V$. In the simplified rotationally-invariant formulation [Dudarev:1998], the extended Hubbard correction reads [Timrov:2021]:
@@
[…]
where $v$ and $\sigma$ label bands and spin, $\mathbf{k}$ runs over the first Brillouin zone (BZ), and $f^{\sigma}_{v,\mathbf{k}}$ are the KS occupations. In Eq. (3), $m\in I$ and $m^{\prime}\in J$, meaning that $m$ runs over the magnetic quantum numbers associated with a fixed principal quantum number $n$ and orbital angular momentum $l$ of atom $I$, and analogously $m^{\prime}$ for atom $J$. The corresponding generalized KS equation can be written as [Campo:2010],
@@
where $\hat{\mathcal{H}}_{\mathrm{KS}}^{\sigma}$ is the KS Hamiltonian of the standard DFT, $\varepsilon_{v,{\bf k}}^{\sigma}$ is the $v$-th band energy at momentum $\bf k$} and spin $\sigma$, and $\hat{\mathcal{V}}_{U+V}^{\sigma}$ is the Hubbard potential that is derived from the Hubbard energy (2) and which reads [Timrov:2021]:
@@
It is important to note that the computed values of $U^{I}$ and $V^{IJ}$ depend on the choice of Hubbard projector functions $\phi^{I}_{m}(\mathbf{r})$ used to define the occupation matrix in Eq. (3) and the Hubbard potential in Eq. (5[Lee:2020, KirchnerHall:2021, Mahajan:2021]. In the following, we briefly describe how Hubbard parameters are defined and computed within LRT and eACBN0. For both methods, we employ the same projector functions, namely atomic orbitals orthogonalized using the Löwdin method [Lowdin:1950, Mayer:2002], to ensure a fair and consistent comparison of the computed $U^{I}$ and $V^{IJ}$ values.
Hubbard parameters from LRT.
Within the LRT framework, Hubbard parameters are defined from the piecewise linearity condition of the total energy [Eq. (1)] with respect to the occupation of the Hubbard manifold and computed as [Cococcioni:2005, Campo:2010]:
@@
DFPT allows this calculation to be performed without supercells, by reformulating LRT in terms of monochromatic reciprocal-space perturbations in the unit cell. This significantly reduces computational cost [Timrov:2018, Timrov:2021]. The response of the occupation matrix is then computed as:
@@
where $\mathbf{q}$ is the wavevector of the monochromatic perturbation, $N_{\mathbf{q}}$ is the total number of perturbations, and $\Delta_{\mathbf{q}}^{s^{\prime}}n^{s\sigma}_{mm^{\prime}}$ is the lattice-periodic response of the occupation matrices to a $\mathbf{q}$-specific perturbation. Here, $I\equiv(l,s)$ and $J\equiv(l^{\prime},s^{\prime})$, with $s$ and $s^{\prime}$ indexing atoms within the unit cell and $l$ and $l^{\prime}$ indexing unit cells; $\mathbf{R}_{l}$ and $\mathbf{R}_{l^{\prime}}$ are the corresponding Bravais lattice vectors. The quantities $\Delta{\mathbf{q}}^{s^{\prime}}n^{s\sigma}_{mm^{\prime}}$ are obtained from the response KS wavefunctions, which are computed by solving $\mathbf{q}$-specific Sternheimer equations. Further details on the DFPT implementation can be found in Refs. Timrov:2018, Timrov:2021. The efficiency of DFPT for calculating $U^{I}$ and $V^{IJ}$ has been demonstrated in numerous studies, with several recent highlights [Gelin:2024, Uhrin:2025, Bastonero:2025, Chang:2025, Binci:2025, dosSantos:2025].
Hubbard parameters from eACBN0.
In the eACBN0 method, the Hubbard parameters are defined by equating the corrective Hubbard energy [Eq. (2)] with the Hartree-Fock interaction energy [Agapito:2015, Lee:2020]. This leads to the following expressions:
@@
where $m\in I$ and $m^{\prime}\in J$, and the summations $\sum_{\{I\}}$ and $\sum_{\{J\}}$ mean that we sum up contributions coming from all equivalent atoms of the same type. For the case of $U^{I}$-only calculations ($I=J$), the second term in Eq. (13) is dropped [Lee:2020], so $\bar{N}^{II,\sigma}_{v,\mathbf{k}}$ reduces to the Mulliken charge of atom $I$ in state $\psi^{\sigma}_{v,\mathbf{k}}(\mathbf{r})$. Moreover, $\bar{N}^{IJ,\sigma}_{v,\mathbf{k}}$ can be viewed as a position-dependent mixing parameter within the Hartree-Fock approximation, encoding on-site and inter-site screening effects in the localized subspace. It is important to note that for a fixed ground-state density, the $U^{I}$ values computed in ACBN0 remain unchanged upon inclusion of inter-site $V^{IJ}$ terms in eACBN0. However, in self-consistent implementations of eACBN0, a coupling between $U^{I}$ and $V^{IJ}$ emerges because the ground state is obtained under the influence of both on-site and inter-site terms in the Hubbard potential (5), with self-consistency enforced for both the electron density and the Hubbard parameters. Further details on the eACBN0 formalism can be found in Refs. Agapito:2015, Lee:2020, TancogneDejean:2020.
@@ RESULTS
Figure 1: Protocol for computing Hubbard parameters within the “self-consistent” scheme using (a) LRT (DFPT) [Timrov:2021] and (b) eACBN0 [Lee:2020]. In the LRT workflow, Hubbard parameters are evaluated self-consistently after the electronic charge density has been fully converged. In contrast, the eACBN0 scheme updates the charge density and the Hubbard parameters simultaneously until convergence is reached. In both workflows, atomic positions are fixed and no structural relaxations are performed.
Hubbard parameters can be computed at different levels of self-consistency. The simplest approach is a “one-shot” calculation, where the parameters are evaluated using the converged DFT charge density while keeping the geometry fixed. A more advanced “self-consistent” approach iteratively updates both the charge density and the corresponding $U$ and $V$ values, but still at a fixed geometry. Finally, in the “structurally self-consistent” scheme, structural relaxations (atomic positions and unit-cell parameters) are included in the self-consistent cycle [Timrov:2021, Bastonero:2025]. In the main text, we report results obtained with the “self-consistent” scheme, i.e. calculations in which the electronic degrees of freedom are converged together with the Hubbard parameters while the geometry remains fixed. In this approach, all $U$ and $V$ values are determined consistently with the self-consistent charge density [Timrov:2021, Lee:2020], as illustrated in Fig. 1. This allows us to isolate purely electronic effects from structural ones and to assess the influence of extended Hubbard interactions in LRT and eACBN0 with high internal consistency. For completeness, in the Supplemental Information (SI) we also present results obtained using the “structurally self-consistent” scheme, in which the self-consistent cycle includes full structural optimization with updated Hubbard parameters and charge density (see Fig. S1 in the SI). In this case, the crystal structure, the Hubbard parameters, and the charge density are iterated to convergence, yielding mutually consistent electronic and crystal structures. All computational details are provided in the Methods section.
@@
Since the band gap is central to this study, it is useful to recall that DFT is not a spectral theory. Local and semilocal xc functionals (LSDA, $\sigma$-GGA) notoriously underestimate band gaps in transition-metal oxides (TMOs) due to strong SIEs, particularly on TM ions. DFT+$U$ can alleviate this issue, but only when the correction is applied to states that define the gap edges (for insulators) or to states near the Fermi level (for spurious DFT metals) [KirchnerHall:2021]. The inter-site $V$ correction can further improve band gaps in covalent systems, but again only when applied to orbitals directly involved in gap formation. Prior work supports these points: Ref. KirchnerHall:2021 showed that DFT+$U$ with LRT-derived $U$ improves agreement with experimental band gaps compared to (semi-)local functionals, while DFT+UU+$V$ with LRT-derived parameters yields even better accuracy in TMOs with strong covalency [Mahajan:2021, Mahajan:2022, Timrov:2023, Gebreyesus:2023]. For $p$-block semiconductors without TM ions, Ref. Lee:2020 demonstrated that DFT+UU+$V$ with eACBN0-derived parameters accurately reproduces experimental band gaps. These insights provide the foundation for our analysis below.
@@
Figure 2: DFT+$U$ and DFT+UU+$V$ self-consistent calculations for TMOs without on-site corrections on O-2pp states ($U_{O}$), using the workflow described in Fig. 1. (a) On-site Hubbard $U_{TM}$ parameter for 3dd orbitals of TM ions and inter-site Hubbard $V_{TM,O}$ parameter between 3dd orbitals of TM ions and 2pp orbitals of oxygen, (b) band gap, and (c) magnetic moment. LRT and eACBN0 denote the method used to obtain the Hubbard parameters. The experimental band gaps and magnetic moments for MnO [Kurmaev2008prb, Elp15301991prb, Cheetham1983prb, Fender1968JCP], FeO [Bowen1975JSSC, Roth1958prb, Battle1979JPCSSP], CoO [Elp60901991prb, Khan1970prb, Jauch2001prb], and NiO [Kurmaev2008prb, Sawatzky1984prl, Cheetham1983prb, Fender1968JCP] are indicated by the dashed blue lines in panels (b) and (c), respectively.
@@
To compare the two methods under as equivalent conditions as possible, we perform LRT and eACBN0 calculations using the “self-consistent” scheme described in Fig. 1 and report the Hubbard parameters in Figs. 2 and 3. We focus first on the $U$ values shown as bar plots (unhatched) in Fig. 2 obtained in the framework of the DFT+$U$ method. A clear trend emerges where $U_{TM}$ increases and magnetic moment decreases with the number of 3dd electrons. Furthermore, both calculation methods do not alter the $U_{TM}$ values significantly when the inter-site interactions are included (compare the hatched and not hatched patterns in Fig. 2(a)). Focusing on MnO, the dominant Mn-$4s$ character of the lowest conduction band turns out to be insensitive to the Hubbard $U$ parameter applied to Mn-$3d$ states. The primary role of $U_{TM}$ is to enhance the splitting between occupied and unoccupied Mn-3dd states. This is confirmed by the partial density of states (PDOS), which shows that with increasing $U_{TM}$, occupied $d$ states shift to lower energies, while the unoccupied $d$ states (lying above the conduction band minimum) shift to higher energies [Yang:2022]. This mechanism is found to be general for other late TMOs like FeO, CoO, and NiO, which also exhibit a prominent TM-4ss character at the conduction band minimum at the $\Gamma$ point. For these materials, we consistently observe that while the energy splitting between occupied and unoccupied $d$ states scales proportionally with $U_{TM}$, the variation in the fundamental band gap remains minor in comparison.
@@
When the inter-site Hubbard interactions ($V_{TM,O}$) between TM 3dd and oxygen 2pp states are included, the differences between the two methods remain almost unchanged. In Fig. 2, the results with the $V_{TM,O}$ term are illustrated with hatched patterns. LRT yields relatively small $V_{TM,O}$ values (below 1 eV), resulting in negligible differences in band gaps and magnetic moments whether $V_{TM,O}$ term is included or not in the total energy calculation. In contrast, eACBN0 consistently produces larger $V_{TM,O}$ values (close to 3 eV) across all TMOs, resulting in more noticeable changes in band gaps compared to LRT, while the magnetic moments change by a small margin. This effect arises because the $V_{TM,O}$ term facilitates electron delocalization into the overlapping regions between atomic pairs. This counteracts the strong localization favored by the $U_{TM}$ term and thus reduces the local magnetic moment, as previously shown through charge density difference analysis in our earlier studies [Yang:2022].